A population has a mean of 100 and equals 20. Given this info, what value would be 1.5 standard deviations above the mean?


1
Expert's answer
2021-02-24T07:01:11-0500

We have that

"\mu = 100"

"\sigma=20"

The value 1.5 standard deviations above the mean:

"\mu +1.5\cdot \sigma = 100+1.5\cdot 20 = 130"


Answer: 130